I fixed some mistakes in the budget:
1. The BS pendulum resonance was corrected from 0.8Hz to 1Hz
2. Added missing X3 filter in the coil filters
3. Optical gain is now computed from MICH to AS55 instead of BS to AS55 and is calculated to be: 9.95e8 cts/m.
4. Coil driver gain is still unmeasured but it is found to be 1.333 to make the actuation calibration from BS to MICH match the measurement (see attachment 1).
Attachment 2 shows the resulting MICH OLTF.
Laser noise was added to the budget in a slightly ad-hoc fashion (will fix later): Yuta and I measured MC_F and computed MC_F*(Schnupp asymmetry)/(Laser frequency). Attachment 3 shows the updated noise budget. |